OpenStudy (anonymous):

Find the exact value of tan (arcsin (2/5)). For full credit, explain your reasoning

OpenStudy (anonymous):

arcsin(2/5)=x then sin x=2/5 tan x = sinx/cosx =sinx/(sqrt(1-sin^2x)) =(2/5)/(sqrt(1-4/5)) or tan(arcsin(2/5)) =(2/5)/(sqrt(21)/5)=2/sqrt(21)

OpenStudy (anonymous):

Thank You! :D